OpenStudy (anonymous):

Solve for x. 8x - 8 = 5x + 16 A. x = 2\frac{1}{3} B. x = 1\frac{11}{24} C. x = 8 D. x = 2

pooja195 (pooja195):

What do you think the first step is?

OpenStudy (anonymous):

Honestly, no clue :/

OpenStudy (anonymous):

My teacher is the best. Barely gives us extra help

pooja195 (pooja195):

We are trying to get x by itself think : How can i get x by itself what operation would i need to use

OpenStudy (mathmale):

Important: combine like terms. Like terms: 2 girls + 3 girls = 5 girls. Unlike terms: 2 girls + 5 boys = 8 boys

OpenStudy (anonymous):

I don't know hwo to combine like terms in Math though.

OpenStudy (anonymous):

8x + 5x = 13?

OpenStudy (anonymous):

13x

pooja195 (pooja195):

Good except thats not what we want to do we want to subtract 8-5=?

OpenStudy (mathmale):

But the 8x is on the left side of your equation and the 5x is on the right. So, we don't add those together. What do we do? Hint: Subtract 5x from both sides of the equation first and then simplify.

OpenStudy (anonymous):

8-5 = 3 and 16 - 5 = 9

OpenStudy (anonymous):

WHOOPS = 11

pooja195 (pooja195):

Why did you subtract 16 and 5?

OpenStudy (anonymous):

I was thinking of the number 9 for something else mb XD

OpenStudy (anonymous):

Subtract 5x from both sides of the equation first and then simplify.

pooja195 (pooja195):

Remember like terms x's with the x's plain numbers with the plain numbers

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

8-5=3 5-5= 0

pooja195 (pooja195):